Write a system of equations and solve. The sum of the measures of the two smaller angles in a triangle is 2 degrees less than the measure of the largest angle. If the middle angle measures 40 degrees less than the largest, find the measures of the angles of the triangle.
I realize that my equations will have an x, y and z for the variable but trying to get those equations is what I am having difficulty with.

Answer by scott8148(6628)

x + y + z = 180

x + y + 2 = z

y + 40 = z